OpenStudy (anonymous):

Which of the following equations represents a horizontal line? A. x = 6 B. y = 2x - 5 C. y = -4

OpenStudy (anonymous):

Write each of the equations in the standard slope - intercept form, y = m*x +b A) -3y = -3x + 6 (Divide each term of the equation by -3) or y = x - 2 B) 3y = - 6x + 6 (Divide each term of the equation by 3) or y = -2x + 2 C) x = -3/2 D) 4y = 6 or y = 3/2 Equation A is a line with slope 1 and y-intercept of -1 (Not a horizontak line) Equation B is a line with slope -2 and y-intercept = 2 Equation C is a VERTICAL line at x = -3/2 Equation D is a hoizontal line at y = 3/2